# Per-antenna power constraints: constructing Pareto-optimal precoders with cubic complexity under non-negligible noise conditions

### v1

@@ I-B Fixed Quality-of-Service convex problem
is nonempty under the power constraints (2). This problem was studied in depth in [11]. Below we will repeat some of the main ideas and results of [11] on the topic, since our contribution will be partially aimed at improving those.
Utilizing $\lvert\alpha\rvert^{2}=\alpha\alpha^{*},\alpha\in\mathbb{C}$, and recalling (1), the set of inequalities (5), (2) can be rewritten as a convex optimization problem [^3] Through the paper, indeces $j$ and $k$ are swapped sometimes, but there is no way to completely avoid that in some derivations with sum reordering.
@@
The trivial objective functional 0 allows the employment of the common convex optimization theory [16], in the form of the Lagrangian and the duality. Assume the Lagrangian multipliers that correspond to quality and power constraints are split into two groups $\hat{\lambda}$ and $\hat{\mu}$. Then, rescaling $\lambda_{j}:=\hat{\lambda}_{j}\gamma_{j}\omega_{j}^{2}$, $\mu_{j}:=\hat{\mu}_{j}\beta_{j}$ gives
$\mathscr{L}(P;\lambda,\mu)=\sum\limits_{j=1}^{m_{ue}}\lambda_{j}-\sum\limits_{i=1}^{m_{tx}}\mu_{i}$ (7)
$+\sum\limits_{j=1}^{m_{ue}}p_{j}^{*}(\begin{bmatrix}\frac{\mu_{1}}{\beta_{1}}&&\\ &\ddots&\\ &&\frac{\mu_{m_{tx}}}{\beta_{m_{tx}}}\\ \end{bmatrix}+\sum\limits_{k\neq j}\frac{\lambda_{k}}{\omega_{k}^{2}}h_{k}h_{k}^{*}-\frac{\lambda_{j}}{\gamma_{j}\omega_{j}^{2}}h_{j}h_{j}^{*})p_{j}.$
By dual problem definition $s(\lambda,\mu):=\inf_{P}\mathscr{L}(P;\lambda,\mu)$ is to be maximized, thus it is only sufficient to consider the case when the matrix in brackets above is nonnegative-definite. Since
$A>0,A-xx^{*}\geq 0\Leftrightarrow x^{*}A^{-1}x\leq 1,$
one can establish
$\gamma_{j}\geq\frac{\lambda_{j}}{\omega_{j}^{2}}h_{j}^{*}(\begin{bmatrix}\frac{\mu_{1}}{\beta_{1}}&&\\ &\ddots&\\ &&\frac{\mu_{m_{tx}}}{\beta_{m_{tx}}}\\ \end{bmatrix}+\sum\limits_{k\neq j}\frac{\lambda_{k}}{\omega_{k}^{2}}h_{k}h_{k}^{*})^{-1}h_{j}.$ (8)
Differentiating Lagrangian (7) and solving $\frac{\partial\mathscr{L}}{\partial p_{j}}=0$ yields the direction of optimal $p_{j}$, and the optimal coefficients can later be found by solving (8) as a set of equalities [11]. Denoting the directions as $\hat{p}_{j}$ and the coefficients as $\kappa_{j}$, that results in
@@
The work [11] (Corollary 3.6) formally proves that any obtainable combination $\gamma_{1},\gamma_{2},\dots,\gamma_{m_{ue}}$ of SINR values is necessarily obtainable by a precoder $P$ built with formulas (9) with some combination of $\mu,\lambda$ with no additional scaling. The formulas (9) are, however, numerically complex: $O(m_{tx}^{3}m_{ue})$ operations are required if they are applied directly. Also, not every combination $\mu,\lambda$ yields a valuable $P$: selecting the multiplier values at random most commonly results in a violation of power constraints. The Pareto-boundary, assuming smoothness, should be a surface that can be described using $m_{ue}-1$ parameters (since it is a surface in $\mathbb{R}^{m_{ue}}$), but the rescaled Lagrange multipliers $\lambda$ and $\mu$ constitute a set of $m_{tx}+m_{ue}$ parameters. One of those can be discarded using
@@
which follows from (7) and the trivial zero-valued optimization functional in (6), but $m_{tx}$ ’extra’ degrees of freedom remain: $m_{tx}$ extra additional equalities should be imposed if we want (9) to become a true Pareto boundary parameterization.
@@ II-A SINR gradients
Assuming $\omega_{k}>0$, $\mathrm{SINR}_{k}$ is a smooth real-valued function of $m_{tx}\times m_{ue}$ complex parameters of $P$; in order to safely use differentiation, let us split the precoding matrix into a real and imaginary part,
$P=\mbox{re}(P)+i\mbox{im}(P);\mbox{re}(P),\mbox{im}(P)\in\mathbb{R}^{m_{tx}\times m_{ue}}.$
Then, by a ’SINR gradient’, we would denote
$\frac{\partial\mathrm{SINR}_{k}}{\partial P}=\frac{\partial\mathrm{SINR}_{k}}{\partial\mbox{re}(P)}+i\frac{\partial\mathrm{SINR}_{k}}{\partial\mbox{im}(P)}.$
@@
the following derivation will be sufficient for our purposes:
  • If
    $\left(\Delta P,\frac{\partial\mathrm{SINR}_{k}}{\partial P}\right)_{F}>0,\mbox{ (and is a real value)}$
    then for all sufficiently small $\epsilon>0$
    $\mathrm{SINR}_{k}(P+\epsilon\Delta P)=\mathrm{SINR}_{k}(P)+\alpha_{k}\epsilon+o(\epsilon),\alpha_{k}>0.$
By direct real-valued differentiation, it follows that $\frac{\partial\mathrm{SINR}_{k}}{\partial P}$ has a closed-form representation as a rank-one matrixx:
A rank-one structure is to be expected, since in the formula for one particular $\mathrm{SINR}_{k}$ each precoding vector is only used in a scalar product with $h_{k}$. As for $d_{k}$, the exact vector values have been carefully checked using numerical derivatives. Since it will appear in further analysis, let us introduce a matrix
@@
$D$ $:=4\begin{bmatrix}\frac{1}{{\cal W}_{1}}&&\\ &\ddots&\\ &&\frac{1}{{\cal W}_{m_{ue}}}\end{bmatrix}G\circ F,$
@@
II-B Looking for a precoder improvement
Let us establish the constraints on the precoding matrix $\Delta P$, sufficient to make the precoding matrix $P+\epsilon\Delta P$ strictly better as compared to $P$ in the Pareto-sense for $\epsilon\rightarrow 0_{+}$. The constraints are straightforward: it would be sufficient to use
$(\Delta P,h_{1}d_{1}^{T})_{F}$ $>0,$ $(\Delta P,e_{2}e_{2}^{*}P)$ $<0,$
$(\Delta P,h_{2}d_{2}^{T})_{F}$ $>0,$ $(\Delta P,e_{3}e_{3}^{*}P)$ $<0,$
$\dots$ $\dots$
$(\Delta P,h_{m_{ue}}d_{ue}^{T})_{F}$ $>0;$ $(\Delta P,e_{m_{tx}}e_{m_{tx}}^{*}P)$ $<0;$
to ensure strict SINR improvements and no power violations. Seeing that the amount of inequalities $m_{tx}+m_{ue}-1$ is much smaller than the number of unknowns $m_{tx}\times m_{ue}$, it would be sufficient to change those to a system of equations, e.g.
$(\Delta P,h_{1}d_{1}^{T})_{F}$ $=1,$ $(\Delta P,e_{2}e_{2}^{*}P)$ $=-1,$
$(\Delta P,h_{2}d_{2}^{T})_{F}$ $=1,$ $(\Delta P,e_{3}e_{3}^{*}P)$ $=-1,$
$\dots$ $\dots$
$(\Delta P,h_{m_{ue}}d_{ue}^{T})_{F}$ $=1;$ $(\Delta P,e_{m_{tx}}e_{m_{tx}}^{*}P)$ $=-1;$
and prove this underdetermined linear system is non-degenerate - then it always has a solution. Vectorizing the unknown $\Delta P$, one could rewrite the system in a matrix form by replacing rank-one matrices with Kronecker products:
All that remains is to show that the long-column matrix $Y\in\mathbb{C}^{m_{tx}m_{ue}\times(m_{tx}+m_{ue}-1)}$ has a full column rank under the conditions of Proposition 1.
II-C Rank of a Hadamard product
@@
Then, we can switch to a Hadamard product:
@@
II-D Rank of an extended SINR derivative matrix
@@ III-A Iterative refinement
From Table I, it is clear that the iteration number growth does not really depend on the sizes of $H,P$, and depends logarithmically on $\frac{1}{\delta}$. In the numerical experiment section we will show that running the algorithm till convergence is mostly interesting for theoretical ’upper-bound’ studies, while for practical application basically one or two computations of $P(\lambda,\mu)$ is sufficient to obtain highly competitive solutions. Moreover, in the next subsection we will argue that the formulas (9) should not be used directly - because there is a way to rewrite those with far better complexity.
@@ III-B Efficient parameterization implementation
Then, recalling (9), note that the values of
$\widehat{P}=\begin{bmatrix}\hat{p}_{1}&\dots&\hat{p}_{m_{ue}}\end{bmatrix}$
and the values of signal and interference coefficients $S$ are defined by two matrices
@@
(Numerical instability) Assuming $C$ is diagonal and $\omega$ approaching 0 in the Algorithm 1 brings us to two divisions of the form
@@
Since $C$ is rescaled by the inverses of $\omega$, $c_{j}$ is a potentially large real value - and it’s clear that the above two operations are dangerous in floating-point arithmetic. Using the Algorithm 1 in double precision with SINR values above $10^{8}$ will lead to $\gamma_{j}$ sometimes popping up as negative because of the rounding errors in $f_{j}\approx 1$. Either avoid these extreme scenarios or use higher precision.
@@
(d) Zero-forcing reference
(e) SLNR reference
Figure 1: Simulation results for the toy channel of size $H\in\mathbb{R}^{8\times 3}$. Each axis represents a value $10\log_{10}(1+\mathrm{SINR}_{k}),k=1,2,3.$ Color shows the mean value. The graphs are constructed pointwise using $1.25\times 10^{7}$ points, and the arrows show the point with the highest mean throughput.
@@ IV-A Reference methods
Let us provide a numerical study of the proposed algorithm framework of combining Algorithm 1 combined with an iterative refinement of Subsection III-A. Since we were aiming at studying the Pareto surfaces of the SINR multi-objective, we would need to build some reference surfaces: let us describe our references. We will use an SVD-notation
@@
    This solution manifests a zero-interference idea $H^{*}P_{\mathrm{ZF}}=I$, but doesn’t satisfy per-antenna power constraints by default; it is commonly post-processed using ’power allocation’, which is essentially column scaling. We will utilize ’power allocation’ to build a surface of ’all possible’ Zero-Forcing solutions by using
@@
    with various weights $\kappa\geq 0$ limited by $\sum\limits_{j=1}^{m_{ue}}\kappa_{j}=1$.
    Surely, we will not be able to check all possible combinations of $\kappa$ since they form a continuum, but we can only build readable surface graphs for $m_{ue}\leq 3$; for this dimensionality a large enough discrete set of vectors $\kappa\in\mathbb{R}_{+}^{m_{ue}}$ will suffice for demonstrations.
    The Zero-Forcing solution is known to be a strong reference when the noise values $\omega_{k}$ are approaching zero: by definition (1) the SINRs provided then approach infinity. However, for non-trivial $\omega_{k}$ the next reference is commonly superior.
@@
    where SLNR stands for ’Signal-to-Leakage-and-Noise-Ratio’ [19, 20]. The replacement is non-equivalent, but SLNR-optimization commonly leads to decent SINR values aswell. Since each $\mathrm{SLNR}_{k}$ only depends on one respective column $p_{k}$ of the precoder, SLNR-optimal directions $p_{k}$ can be computed explicitly (resulting in (29)), but $P_{\mathrm{SLNR}}$ still doesn’t take the per-antenna power constraints into account. Then again, we will utilize ’power allocated’
@@
    as a reference surface parameterized by nonnegative $\kappa$ that sum into one. $P_{\mathrm{SLNR}}$ maintains a complexity of $O(m_{tx}m_{ue}^{2})$ and automatically collapses into $P_{\mathrm{ZF}}$ when $\omega\rightarrow 0$, but commonly offers superior SINR values for nontrivial $\omega$. Both ZF and SLNR still only use the column space $U$ of the input channel $H$ - which is sufficient for the Frobenius total power constraint (3), but is strictly suboptimal experimentally under the per-antenna power constaints (2), which are the focus of this paper.
  • Quality-of-Service convex solver ([11]). Since the fixed Quality-of-Service problem (5) is proved to be convex, it can, with some modifications, be solved using convex solvers such as the interior-point method. The work [11] provides explanation how the SINR constraints can be rewritten as second-order cone constraints, and it is possible to introduce a relaxation parameter such that the starter interior point can be constructed explicitly, but we will omit the unnecessary details and refer to [21, 11, 22] for more information on QoS convex solvers. The important outline is
    • The Quality-of-Service problem (5) constraints can be rewritten as second-order cone constraints [11]: it can be solved using a convex solver for each positive combination of $\gamma_{1}\dots\gamma_{m_{ue}}$.
    • Assuming that a whole precoder $P$ is an unknown in the convex solver, a direct implementation of a convex solver results in an iterative algorithm with an iteration complexity of $O(m_{tx}^{3}m_{ue}^{3})$ (cubic in the number of unknowns), which can be reduced down to fourth-order complexity using some inherent matrix structures - but is still high.
    For the purpose of building an ideal reference, we can select a ray direction vector $\gamma\in\mathbb{R}^{m_{ue}}_{+}$, and accurately intersect the ray with a Pareto boundary using line-search along one positive scaling constant - which will give us one correct Pareto-boundary point. This reference is very slow and unusable in real-time, but we will use it to verify that our developed parametric precoding scheme arrives at the true Pareto boundary.
@@
IV-B 3D experiments on a toy channel
Let us begin the presentation of our numerical experiments. We will begin with a toy real-valued channel of the form
$H=\begin{bmatrix}-0.4&-0.2&0.0\\ -0.4&0.9&1.2\\ -0.2&0.1&-0.2\\ 0.7&-0.8&-0.5\\ 0.0&-0.2&-0.2\\ 0.6&-0.5&0.6\\ -0.4&1.2&-0.1\\ -1.2&-0.4&-0.8\end{bmatrix}\in\mathbb{R}^{8\times 3}.$ (31)
with the noise values and per-row power limits equal to one
$\omega_{k}=1,\beta_{i}=1.\mbox{ }(k\in\{1,\dots,m_{ue}\};\mbox{ }i\in\{1,\dots,m_{tx}\})$
to avoid any type of confusion with squares. This example is considered because it is easily reproducible, allows making explicit 3D Pareto-boundary graphs and helps verifying the previous theoretical developments - if the reader is only interested in simulations with complex channels and higher user numbers, we refer to Subsection IV-E and further.
The system throughput, which is a logarithmic function of SINR (4) is sometimes weighted differently in applications. The ’per-user throughput’ values we will be presenting on the graphs will be
$10\log_{10}(1+\mathrm{SINR}_{k}),$ (32)
which is $1+\mathrm{SINR}$ in decibels’ - changing the logarithm base corresponds to simple scaling if necessary. We will provide some direct SINR values aswell to avoid any confusion.
The Figure 1 represents the surfaces obtained by the proposed algorithm and the two real-time applicable references of ZF and SLNR on the toy real channel. In order to show that a large number of iterations is not really necessary for the proposed algorithm in applications - we are giving three surfaces:
Each dimension of the graphs on Figure (1) gives a value of (32) for one of the three users - and the color represents the averaging of those, which is a scaled value of total throughput. It can be seen that running even one $\mu$ update iteration gives a quite competitive result, hence the number of iterations used can be leveraged to balance between optimality and complexity easily in applications.
Each of the surfaces provided is plotted pointwise, by using $1.25\times 10^{7}$ points. The arrow on each graph shows a discrete point at which the highest value of the mean throughput was encountered. Since the granularity is quite low, this value should be an accurate estimate of the highest mean throughput achievable by the corresponding precoder type. Tables III, IV in the Appendix give detailed informations about these points - in case the reader wants to reproduce the results.
TABLE II: Success chance of improving each SINR value a developed precoding algorithm with tolerance $\delta$ by a factor $\upsilon$, 100 runs
$P\in\mathbb{C}^{24\times 8},\delta=10^{-4}$ 0 % 6 % 47 % 82 % 100 %
$P\in\mathbb{C}^{192\times 24},\delta=10^{-4}$ 0 % 9 % 28 % 44 % 95 %

### v2

@@ I-B Fixed Quality-of-Service convex problem
is nonempty under the power constraints (2). This problem was studied in depth in [11]: utilizing $\lvert\alpha\rvert^{2}=\alpha\alpha^{*},\alpha\in\mathbb{C}$, and recalling (1), the set of inequalities (5), (2) can be rewritten as a convex optimization problem
@@
By analyzing the corresponding Lagrangian and its derivatives, [11] (Corollary 3.6) formally proves that any obtainable combination $\gamma_{1},\gamma_{2},\dots,\gamma_{m_{ue}}$ of SINR values is necessarily obtainable by a precoder $P$ built with formulas
@@
with some combination of Lagrange multipliers $\mu,\lambda$ with no additional scaling. The formulas (7) are, however, numerically complex: $O(m_{tx}^{3}m_{ue})$ operations are required if they are applied directly. Also, not every combination $\mu,\lambda$ yields a valuable $P$: selecting the multiplier values at random most commonly results in a violation of power constraints. The Pareto-boundary, assuming smoothness, should be a surface that can be described using $m_{ue}-1$ parameters, but the Lagrange multipliers $\lambda$ and $\mu$ constitute a set of $m_{tx}+m_{ue}$ parameters. One of those can be discarded using
@@
which is proved in [11] by recalling that the optimization functional in (6) is trivial, but $m_{tx}$ ’extra’ degrees of freedom remain: $m_{tx}$ additional equations should be imposed if we want (7) to become a true Pareto boundary parameterization.
@@ II-A SINR gradients
Assuming $\omega_{k}>0$, $\mathrm{SINR}_{k}$ is a smooth real-valued function of $m_{tx}\times m_{ue}$ complex parameters of $P$; in order to safely use differentiation, let us split the precoding matrix into a real and imaginary part, and denote
$\frac{\partial\mathrm{SINR}_{k}}{\partial P}=\frac{\partial\mathrm{SINR}_{k}}{\partial\mbox{re}(P)}+i\frac{\partial\mathrm{SINR}_{k}}{\partial\mbox{im}(P)}$
as the ’SINR gradient’. By direct real-valued differentiation, it follows that $\frac{\partial\mathrm{SINR}_{k}}{\partial P}$ has a closed-form representation as a rank-one matrixx:
A rank-one structure is to be expected, since in the formula for one particular $\mathrm{SINR}_{k}$ each precoding vector is only used in a scalar product with $h_{k}$. As for $d_{k}$, the exact vector values have been carefully checked using numerical derivatives.
@@
and showing that a $\Delta P\in\mathbb{C}^{m_{tx}\times m_{ue}}$ exists that suffices the system of equations
$(\Delta P,h_{1}d_{1}^{T})_{F}$ $=1,$ $(\Delta P,e_{2}e_{2}^{*}P)_{F}$ $=-1,$
$(\Delta P,h_{2}d_{2}^{T})_{F}$ $=1,$ $(\Delta P,e_{3}e_{3}^{*}P)_{F}$ $=-1,$
$\dots$ $\dots$
$(\Delta P,h_{m_{ue}}d_{ue}^{T})_{F}$ $=1;$ $(\Delta P,e_{m_{tx}}e_{m_{tx}}^{*}P)_{F}$ $=-1;$
would be sufficient [^3] The equations enforce that the direction $\Delta P$ has a positive cosine with SINR derivatives and a negative cosine with power derivatives. They are sufficient but not necessary, of course: at the very least, the values of $\pm 1$ can be replaced with arbitrary positive and negative values respectively. to show that the precoding matrix $P+\epsilon\Delta P$ strictly better as compared to $P$ in the Pareto-sense for $\epsilon\rightarrow 0_{+}$. Vectorizing the unknown $\Delta P$, one could rewrite the system in a matrix form by replacing rank-one matrices with Kronecker products:
Consequently, all that remains is to show that the long-column matrix $Y\in\mathbb{C}^{m_{tx}m_{ue}\times(m_{tx}+m_{ue}-1)}$ has a full column rank under the conditions of Proposition 1. Since it will be useful in the further analysis, let us introduce a matrix notation
@@
$D$ $=4\begin{bmatrix}\frac{1}{{\cal W}_{1}}&&\\ &\ddots&\\ &&\frac{1}{{\cal W}_{m_{ue}}}\end{bmatrix}G\circ F,$
@@
II-B Rank of a Hadamard product
@@
Then, we can switch to analyzing Hadamard product by using
@@
II-C Rank of an extended SINR derivative matrix
@@ III-A Iterative refinement
From Table I, it is clear that the iteration number growth does not really depend on the sizes of $H,P$, and depends logarithmically on $\frac{1}{\delta}$. In the next subsection we will argue that the formulas (7) should not be used directly - because there is a way to rewrite those with far better complexity.
@@ III-B Efficient parameterization implementation
Then, recalling (7), note that the values of $\hat{p}_{j}$ and the values of signal and interference coefficients $S$ are defined by two matrices
@@
(Numerical instability) Assuming $C$ is diagonal and $\omega\rightarrow 0_{+}$ in the Algorithm 1 brings us to two divisions
@@
Since $C$ is rescaled by the inverses of $\omega$, $c_{j}$ is a potentially large real value - and the above two operations are dangerous in double precision with SINR values above $10^{8}$. To prevent $\gamma_{j}$ showing up as negative because of the rounding errors in $f_{j}\approx 1$, either avoid extreme scenarios or use higher precision.
@@
$H=\begin{bmatrix}-0.4&-0.2&0.0\\ -0.4&0.9&1.2\\ -0.2&0.1&-0.2\\ 0.7&-0.8&-0.5\\ 0.0&-0.2&-0.2\\ 0.6&-0.5&0.6\\ -0.4&1.2&-0.1\\ -1.2&-0.4&-0.8\end{bmatrix}\in\mathbb{R}^{8\times 3}.$
$\omega_{k}=1,\beta_{i}=1.$
$k\in\{1,\dots,m_{ue}\};\mbox{ }i\in\{1,\dots,m_{tx}\}$
(d) Channel values, problem parameters and ZF/SLNR references
Figure 1: Simulation results on a toy channel of size $H\in\mathbb{R}^{8\times 3}$. Each axis represents a value $10\log_{10}(1+\mathrm{SINR}_{k}),k=1,2,3.$ Color shows the mean value. The graphs are constructed pointwise using $1.25\times 10^{7}$ points, and the arrows show the point with the highest mean throughput.
@@ IV-A Reference methods
Let us provide a numerical study of the proposed algorithm framework of combining Algorithm 1 combined with an iterative refinement of Subsection III-A. Since we were aiming at studying the Pareto surfaces of the SINR multi-objective. Let us describe our references. We will use an SVD-notation
@@
    This solution manifests a zero-interference idea $H^{*}P_{\mathrm{ZF}}=I$, and is commonly post-processed by
@@
    with various weights [^7] Surely, we will not be able to check all possible ’power allocations’ of $\kappa$ since they form a continuum, but we can only build readable surface graphs for $m_{ue}\leq 3$; for this dimensionality a large enough discrete set of vectors $\kappa\in\mathbb{R}_{+}^{m_{ue}}$ will suffice for demonstrations. $\kappa\geq 0$ limited by $\sum\limits_{j=1}^{m_{ue}}\kappa_{j}=1$.
@@
    where SLNR stands for ’Signal-to-Leakage-and-Noise-Ratio’ [18, 19]. Again, we will utilize ’power allocated’
@@
    as a reference surface parameterized by nonnegative $\kappa$ that sum into one. $P_{\mathrm{SLNR}}$ maintains a complexity of $O(m_{tx}m_{ue}^{2})$ and collapses into $P_{\mathrm{ZF}}$ when $\omega\rightarrow 0$, but commonly offers superior SINR values for nontrivial $\omega$.
  • Quality-of-Service convex solver. ([20, 11, 21, 22]) For the purpose of building an ideal reference, we can select a ray direction vector $\gamma\in\mathbb{R}^{m_{ue}}_{+}$, and accurately intersect the ray with a Pareto boundary using a QoS solver of (6) and a line-search along one positive scaling constant - which will give us one correct Pareto-boundary point. This reference is very slow [^8] One iteration of an interior point method for (6) has a naive complexity of $O(m_{tx}^{3}m_{ue}^{3})$, reducible down to fourth order complexity using the innate matrix structures. and unusable in real-time, but we will use it to verify that our developed parametric precoding scheme accurately arrives at the true Pareto boundary.
@@
IV-B 3D experiments on a toy channel
We will begin our numerical experiments with a toy fixed channel $H\in\mathbb{R}^{8\times 3}$. This example is considered because it is easily reproducible and allows making explicit 3D Pareto-boundary graphs - if the reader is only interested in simulations with complex channels and higher user numbers, we refer to Subsection IV-E and further.
The Figure 1 provides the considered parameters and the surfaces obtained by the proposed algorithm and the two references of ZF and SLNR. In order to show that a large number of iterations is not really necessary for the proposed algorithm in applications - we are providing three surfaces:
It can be seen that running even one $\mu$ update iteration gives a quite competitive result, hence the number of iterations used can be leveraged to balance between optimality and complexity easily in applications. Each of the surfaces provided is plotted pointwise, by using $1.25\times 10^{7}$ points. The arrow on each graph shows a discrete point at which the highest value of the mean throughput was encountered. If the reader wants to reproduce the results, we are providing detailed precoding matrix values for these points in Supplementary materials.